Starting from general equilibrium, what would be the long-run effects of a simultaneous reduction in government purchases (G↓) and increase in the money supply (M↑) designed to leave real GDP the same on each of the following economic variables? For each, you should write one of the following responses:  Up (U), Down (D), orSame (S)


  1. The real interest rate (r)
  2. Investment (I)
  3. Consumption (C)
  4. The price level (P)
  5. Budget deficit (G – T)
  6. Future standard of living (i.e., future per capita consumption)
1
Expert's answer
2021-05-05T16:41:29-0400

We can take a standard IS-LM framework to understand the economic activities. 

In an equilibrium, the commodity and the money market should look like the following, (symbols bear usual significance).

"Y=C(Y-T)+I(r)+G"

"\frac{M}{P}=L(r,Y)"


we consider a closed economy.

Now, G falls and M rises. Y remains the same.

Now, we look at the effect on different variables. 


1. Real Interest Rate (r): 

if G falls, that means that releases tax burden on citizens and that raises savings. Savings is the supply of loanable funds. Hence, loans become cheaper, i.e. interest rate falls. 

On the other hand, if money supply rises, from the money market, interest rate falls. 

So, given that price is constant, Real Interest Rate goes DOWN.


2. Investment (I)

With falling interest rate, investment rises. That makes sense, because if Y remains unchanged, and G falls, the fall in G has to be compensated. 

So, Investment goes UP.


3. Consumption: 

If M rises, that means there is more money in the economy to be spent on. That increase private consumption. 

With falling G, tax burden is relaxed, if the MPC is high, then excess disposable income is consumed. 

So, consumption goes UP


4. The Price Level.

The reduction in government spending means a contraction in aggregate demand. So, there will be an excess supply and price level falls. 

On the other hand, with expansionary monetary policy, inflation takes place. Hence price level rises. 

Now, since there cannot be any excess supply and full employment output is maintained. We can safely say that price level remains unchanged. 

So, The Price Level remains SAME.


5. Budget Deficit:

If G falls then (G - T) falls obviously. 

On the other hand, if M rises, a part of budget deficit might be financed by newly printed money. 

So, in totality, budget deficit falls. 

Budget deficit goes DOWN.


Answers

The real interest rate (r) goes DOWN.

Investment (I) goes UP.

Consumption (C) goes UP.

The price level (P) goes SAME.

Budget deficit (G – T) goes DOWN.